The solution to" the product of 5 and the quantity decreased by 10 is 40" ?
Answer by ankor@dixie-net.com(22740)

The solution to" the product of 5 and the quantity decreased by 10 is 40" ?
:
Let x = "the quantity"
then
5(x-10) = 40
5x - 50 = 40
5x = 40 + 50
5x = 90
x = 90/5
x = 18
:
:
See if that checks out: 5(18-10) = 40